If 200 labor hours were required to produce the first unit in a production run, and n labor hours were required to produce the seventh unit. Find the value of n if the learning curve rate during production is 80%.?
Answer #1

Ans) TN = T1 (Nb)

Where, TN is time for nth unit

T1 is time required for 1st unit.

b = (log of learning rate)/(log2)

TN = 200 (7(log0.8/log2)) = 200 × 7(-0.3219) = 200 × 0.5345 =106.8979